how do you solve this problem
y=3x
5x+2y=44

Answer by checkley77(12844)

y=3x
5x+2y=44 replace y with (3x) in this equation & solve for x.
5x+2(3x)=44
5x+6x=44
11x=44
x=44/11
x=4 ans.
y=3*4
y=12 ans.
Proof:
5*4+2*12=44
20+24=44
44=44